# Towards Fuzzing Zero-Knowledge Proof Circuits

### v1

@@ 5. Related Work
ZKP Security and Automated Analysis. Recent studies have categorized ZKP bugs and threats to different layers, highlighting a comprehensive taxonomy of vulnerabilities (Chaliasos et al., 2024a; Tang et al., 2024). Several works have also focused on attacking ZKP constructions, demonstrating how oversights can lead to severe security vulnerabilities in all ZKP implementations that use the specific constructions (Dao et al., 2023; Khovratovich et al., 2025; Ciobotaru et al., 2024). Another line of work has focused on rigorous verification of ZKP components, for instance, verifying the verifier itself to ensure correctness (Firsov and Livshits, 2024). While formal verification has been used to prove the absence of bugs in ZKP circuits (Pailoor et al., 2023; Liu et al., 2023; Coglio et al., 2023b, a; Ozdemir et al., 2023; Isabel et al., 2024; Stronati et al., 2024), it often faces scalability challenges. Similarly, static analysis approaches (Wen et al., 2023; Soureshjani et al., 2023) can detect certain classes of issues but may suffer from precision limitations and heuristics that miss broader categories of bugs. These hurdles underscore the need for complementary, scalable, and precise techniques, such as fuzzing, that could reveal critical bugs.
Fuzzing ZKP Implementations. Fuzzing ZKP implementations has gained traction in a few domain-specific contexts. For instance, recent work identified completeness bugs in a specific zkEVM implementation by systematically mutating and monitoring execution paths (Li et al., 2024). Further, metamorphic testing has been employed to uncover errors in ZKP compilers (Hochrainer et al., 2024; Xiao et al., 2025). Finally, SnarkProbe (Fan et al., 2024) focuses on revealing bugs in ZKP DSL backends. These papers either focus on non-circuit bugs or struggle with the challenges we outlined in Section 3, such as solving the oracle problem for detecting soundness bugs.
@@ 6. Conclusion
We have discussed the unique challenges of fuzzing ZKP circuits, including issues with test oracles, input/witness generation, and performance overheads, and proposed practical solutions such as differential testing, invariant-based oracles, and hybrid methods to enhance bug detection. Our zk-regex case study demonstrates that even a simple fuzzing campaign can reveal significant bugs, highlighting the potential of fuzzing to improve ZKP security. The remaining open problems we have identified provide fruitful ideas for future fuzzing research in this domain.

### v2

@@ 5. Related Work
ZKP Security and Automated Analysis. Recent studies have categorized ZKP bugs and threats to different layers, highlighting a comprehensive taxonomy of vulnerabilities (Chaliasos et al., 2024a). Several works have also focused on attacking ZKP constructions, demonstrating how oversights can lead to severe security vulnerabilities in all ZKP implementations that use the specific constructions (Dao et al., 2023; Khovratovich et al., 2025; Ciobotaru et al., 2024). Another line of work has focused on rigorous verification of ZKP components, for instance, verifying the verifier itself to ensure correctness (Firsov and Livshits, 2024). While formal verification has been used to prove the absence of bugs in ZKP circuits (Pailoor et al., 2023; Coglio et al., 2023b; Ozdemir et al., 2023; Isabel et al., 2024), it often faces scalability challenges. Similarly, static analysis approaches (Wen et al., 2024; Soureshjani et al., 2023) can detect certain classes of issues but may suffer from precision limitations and heuristics that miss broader categories of bugs. These hurdles underscore the need for complementary, scalable, and precise techniques, such as fuzzing, that could reveal critical bugs.
Fuzzing ZKP Implementations. Fuzzing ZKP implementations has gained traction in a few domain-specific contexts. For instance, recent work identified completeness bugs in a specific zkEVM implementation by systematically mutating and monitoring execution paths (Li et al., 2024). Further, metamorphic testing has been employed to uncover errors in ZKP compilers (Hochrainer et al., 2024; Xiao et al., 2025). Finally, SnarkProbe (Fan et al., 2024) focuses on revealing bugs in ZKP DSL backends. These papers either focus on non-circuit bugs or struggle with the challenges we outlined in Section 3, such as solving the oracle problem for soundness bugs.
@@ 6. Conclusion
We have discussed the unique challenges of fuzzing ZKP circuits, including issues with test oracles, input/witness generation, and performance overheads, and sketched practical solutions such as differential testing, invariant-based oracles, and hybrid methods to enhance bug detection. Our zk-regex case study demonstrates that even a simple fuzzing campaign can reveal significant bugs, highlighting the potential of fuzzing to improve ZKP security. The remaining open problems we have identified provide fruitful ideas for future fuzzing research in this domain.
Acknowledgements.
We thank the anonymous reviewers for their constructive feedback. Further, we would like to thank Shreyas Londhe and Aayush Gupta from zkEmail for firmly confirming our bugs and brainstorming potential extensions of our fuzzing framework for zk-regex. We also want to thank Kostas Ferles, Burak Kadron, and Ben Mariano from Veridise Inc., for fruitful discussions on hybrid fuzzing techniques for ZK circuits. This work has been partially funded by Aztec Labs.